# DEER: A Benchmark for Evaluating Deep Research Agents on Expert Report Generation

### v2

@@ Appendix D Foundations of Deep Research Report Evaluation Taxonomy
Sensitive Handling corresponds to ethical-screening requirements that evaluate whether sensitive domains are addressed with neutrality, respect, and responsible language in accordance with established editorial and reporting ethics standards International Committee of Medical Journal Editors (2025); Bastian and Moher (2021); Oral History Association (2024); Organization of American Historians (2018); British Psychological Society (2021); National Commission for the Protection of Human Subjects of Biomedical and Behavioral Research (1979); American Political Science Association (2012a). Safety & Impact aligns with harm-assessment provisions that examine whether potential adverse effects, misuse risks, or disproportionate impacts are identified and mitigated under recognized guidelines for responsible scientific communication International Committee of Medical Journal Editors (2025); Guyatt et al. (2013); National Research Council (2011); World Medical Association (2013); Ecological Society of America (2021). Perspective Balance reflects fairness-evaluation criteria that assess whether multiple viewpoints and counterpositions are represented without undue bias following principles of impartiality articulated in authoritative publication and reporting frameworks International Committee of Medical Journal Editors (2025); Bastian and Moher (2021); British Philosophical Association (2024); Australasian Association of Philosophy (2023); CFA Institute (2024); American Bar Association (2023); American Association for Public Opinion Research (2021).
@@
Evidence Coverage aligns with evidence-provision requirements that evaluate whether claims are accompanied by verifiable supporting sources in accordance with established evidentiary and reporting guidelines Bastian and Moher (2021); Data Citation Synthesis Group (2014); Chang et al. (2024); International Society for Stem Cell Research (2025); CFA Institute (2020). Claim Amount corresponds to content-adequacy criteria that assess whether a report supplies all necessary factual and contextual material needed to justify its reasoning under recognized completeness and transparency standards Bastian and Moher (2021); International Organization for Standardization (2011); EQUATOR Network (2025); Linguistic Society of America (2024); The Journal of Organic Chemistry (2025); U.S. Geological Survey (2024); Global Reporting Initiative (2023). Citation Amount align with source-attribution rules that examine whether in-text citations are provided at appropriate argumentative locations following authoritative norms for evidentiary traceability International Committee of Medical Journal Editors (2025); Wilkinson et al. (2016); IEEE (2025); American Chemical Society (2021); Harvard Law Review Association (2020). Reference Amount are evaluated to determine whether the number of actually used references is appropriate and whether they collectively form a well-documented, traceable, and accessible evidence base Stanford HAI (2025); Data Citation Synthesis Group (2014); American Political Science Association (2012b).
@@
Claim Factuality reflects evidence-verification provisions that assess whether factual assertions in a report correspond to verifiable sources and documented evidence traces within established evaluative frameworks Wilkinson et al. (2016); American Physical Society (2023); National Institute of Standards and Technology (2023); Navas and others (2024); National Institute of Standards and Technology (2024); U.S. Securities and Exchange Commission (2024). Citation Support corresponds to source-justification checks that evaluate whether cited references substantively support the claims they accompany according to recognized standards for evidential accountability Wilkinson et al. (2016); American Sociological Association (2018); American Historical Association (2024); Berez-Kroeker and others (2018). Reference Support aligns with source-quality and provenance requirements that examine whether referenced materials reliably and transparently support the arguments for which they are cited, in line with authoritative data-governance and reporting criteria Wilkinson et al. (2016); Chang et al. (2024); Data Citation Synthesis Group (2014); American Psychological Association (2025); International Accounting Standards Board (2024). Reference Quality evaluates whether the set of sources consists of reproducible, trustworthy, and methodologically sound information sources, ensuring that the report’s evidentiary foundation is grounded in reliable materials American Chemical Society (2021); IEEE (2025); Stanford Encyclopedia of Philosophy (2025); World Medical Association (2013); National Commission for the Protection of Human Subjects of Biomedical and Behavioral Research (1979). Reference Diversity evaluates whether the evidence base avoids excessive concentration on a small number of sources and instead maintains a balanced level of source diversity, reducing the risk of narrow, biased, or selectively framed arguments EQUATOR Network (2025); Linguistic Society of America (2024); American Association for Public Opinion Research (2021).
To establish a robust and universally applicable Deep Research Report Evaluation Taxonomy, this framework comprehensively integrates authoritative standards and guidelines from a wide spectrum of disciplines, ranging from quantitative sciences and engineering to the humanities, social sciences, and professional fields such as law and finance (see Table 10). By synthesizing the core principles shared across these diverse domains, this taxonomy prioritizes criteria that are universally recognized as essential for high-quality intellectual work. Consequently, the resulting evaluation model ensures that the generated reports not only adhere to domain-specific best practices but also satisfy the fundamental requirements of validity, clarity, and professional integrity demanded by the global research and practitioner communities.
@@
Appendix E Rubric Structure
In this appendix, we summarize the hierarchical structure and scoring method of the rubric used to evaluate the Deep Research report.
@@ Appendix E Rubric Structure
2. Analytical Soundness └ 2.2 Reasoning 2.2.5 Criterion All major claims follow logically from the previously presented facts, data, interpretations, and assumptions, without skipped steps or unsupported leaps. [2.2.5.1 Coverage] All claims that require logical support are explicitly linked to the relevant facts, data, and interpretations (including the key evidence specified in the Expert Evaluation Guidance, EG), and no central claim is left without the necessary supporting evidence. [2.2.5.2 Quality] Each major claim is explained through clear, well-structured reasoning that shows how the underlying facts, data, and interpretations lead to that conclusion, making the logical connection explicit and easy for an expert reader to follow.
Table 11: Example Coverage/Quality rubric factors for a reasoning-focused element under 2. Analytical Soundness / 2.2 Reasoning.
E.1 Hierarchical Levels
We have the following 4-level hierarchical structure to systematically evaluate report quality in multiple dimensions.
@@ E.1 Hierarchical Levels
Detailed criteria that decompose each dimension into finer evaluation units; 2–5 sub-dimensions are defined under each dimension. This forms an intermediate-level unit that specifies which aspects to examine for each dimension.
Level 3: Evaluation Elements (46)
Evaluation elements that concretize each detailed criterion into units actually observable in the report. One or more elements are mapped under one detailed criterion, and are defined as evaluation units directly corresponding to report content, such as "Inclusion of requested items," "Specification of scope/limitations," and "Specificity of final message."
Level 4: Rubric Factors (101)
Rubric factors corresponding to the fourth level of the rubric are defined to evaluate each evaluation element (Level 3) from two perspectives. One element is subdivided into multiple factors from the perspectives of Coverage (C) and Quality (Q), and these factors serve as the minimum unit for actual scoring.
Thus, with 7 dimensions, 25 sub-dimensions, and 46 elements as the framework, each element is divided into Coverage/Quality factors to constitute a total of 101 specific check items. Table 11 shows an example of one element and its corresponding Coverage/Quality factors in the 2. Analytical Soundness / 2.2 Reasoning dimension.
E.2 Coverage and Quality Factors
The rubric factors corresponding to Level 4 of the rubric evaluate each evaluation element (Level 3) from two perspectives: Coverage (C) and Quality (Q). The Coverage factor examines whether the element has been covered without omission as required throughout the document, i.e., how faithfully it includes content scattered in multiple locations or composed of multiple detailed requirements. On the other hand, the Quality factor evaluates whether the description possesses sufficient depth, logic, and rigor based on professional report standards, assuming the parts already written, i.e., “how well it is written.” By separating Coverage and Quality in this way, we can independently measure (1) what is missing and (2) what the level of the included content is in a long report. In our rubric, Level 4 consists of 101 factor-level items in total, comprising 66 Coverage items and 35 Quality items. Both factors use a 1–10 point scale, and sub-dimensions for scoring are presented in Table 12.
@@ E.2 Coverage and Quality Factors
E.3 Scoring and Aggregation
The score aggregation flow in this evaluation framework consists of factor $\rightarrow$ element $\rightarrow$ criterion $\rightarrow$ dimension. First, rubric factors (the lowest items) are evaluated with a 1–10 score (or N/A). Coverage (C) and Quality (Q) are calculated separately up to the factor/element level, then integrated at the element level, and the resulting element score is raised to the criterion and dimension levels. N/A items are excluded from all average calculations.
Let $r$ be the report, and type $T\in\{C,Q\}$ represent Coverage and Quality, respectively. Let the score of each rubric factor $f$ be $s_{r,f}\in\{1,\dots,10\}$ (or N/A), and the set of non-N/A factors of type $T$ belonging to element $e$ be $F_{e,r}^{T}$. Let $E_{c,r}$ be the set of elements belonging to criterion $c$, and $\mathcal{C}_{d,r}$ be the set of criteria belonging to dimension $d$. Then, the factor $\rightarrow$ element $\rightarrow$ criterion $\rightarrow$ dimension aggregation is defined as follows:
The Coverage/Quality score (element level) of report $r$ for element $e$ is
$S_{r}^{T}(e)=\frac{1}{\lvert F_{e,r}^{T}\rvert}\sum_{f\in F_{e,r}^{T}}s_{r,f},\quad T\in\{C,Q\},$ (3)
and if $F_{e,r}^{T}=\varnothing$, $S_{r}^{T}(e)$ is set to N/A.
The integrated element score $S_{r}(e)$ of element $e$ is set as the average of defined values among C/Q. Let the set of types defined in element $e$ be $\mathcal{T}_{e}=\{T\in\{C,Q\}\mid S_{r}^{T}(e)\neq\text{N/A}\}$.
$S_{r}(e)=\frac{1}{\lvert\mathcal{T}_{e}\rvert}\sum_{T\in\mathcal{T}_{e}}S_{r}^{T}(e),$ (4)
If $\mathcal{T}_{e}=\varnothing$, $S_{r}(e)$ is set to N/A.
The score of criterion $c$ is defined as the average of element scores belonging to that criterion.
$S_{r}(c)=\frac{1}{\lvert E_{c,r}\rvert}\sum_{e\in E_{c,r}}S_{r}(e).$ (5)
If $E_{c,r}=\varnothing$, $S_{r}(c)$ is set to N/A.
The score of dimension $d$ is defined as the average of criterion scores belonging to that dimension.
$S_{r}(d)=\frac{1}{\lvert\mathcal{C}_{d,r}\rvert}\sum_{c\in\mathcal{C}_{d,r}}S_{r}(c).$ (6)
If $\mathcal{C}_{d,r}=\varnothing$, $S_{r}(d)$ is set to N/A.
In summary, we calculate $S_{r}^{C}(e),S_{r}^{Q}(e)$ by averaging C/Q factors in each element, integrate them to create the element score $S_{r}(e)$, then average the element scores by criterion to calculate $S_{r}(c)$, and finally average the criterion scores to calculate the dimension score $S_{r}(d)$.
Appendix F Human-based Information Verification Protocol
F.1 Overview
@@
F.2 Step 1: Claim Extraction and Classification
@@
Evaluators segmented the report into sentences (in ‘Lx.Sy‘ format), reviewed each sentence individually to identify sentences containing claims, and extracted only the core claims. Pronouns were replaced with explicit references according to the context, and if a single sentence contained multiple claims, they were separated. All claims were classified into 6 types (A–F): Explicit Citation (A), Implicit – Same Section (B), Implicit – Previous Section (C), Structural Recap (D), No Citation Required (E), and Unknown Source (F). For A–C type claims, the corresponding citation or evidence position was recorded together. This process was performed on 2 randomly selected reports (total 728 claims), thereby constructing a Ground Truth for evaluating the recall of the extraction model.
@@
F.3 Step 2: Claim Verification

### v3

@@ C.2 Construction Procedure and Evidence Mapping
Sensitive Handling corresponds to ethical-screening requirements that evaluate whether sensitive domains are addressed with neutrality, respect, and responsible language in accordance with established editorial and reporting ethics standards (International Committee of Medical Journal Editors, 2025; Bastian and Moher, 2021; Oral History Association, 2024; Organization of American Historians, 2018; British Psychological Society, 2021; National Commission for the Protection of Human Subjects of Biomedical and Behavioral Research, 1979; American Political Science Association, 2012a). Safety & Impact aligns with harm-assessment provisions that examine whether potential adverse effects, misuse risks, or disproportionate impacts are identified and mitigated under recognized guidelines for responsible scientific communication (International Committee of Medical Journal Editors, 2025; Guyatt et al., 2013; National Research Council, 2011; World Medical Association, 2013; Ecological Society of America, 2021). Perspective Balance reflects fairness-evaluation criteria that assess whether multiple viewpoints and counterpositions are represented without undue bias following principles of impartiality articulated in authoritative publication and reporting frameworks (International Committee of Medical Journal Editors, 2025; Bastian and Moher, 2021; British Philosophical Association, 2024; Australasian Association of Philosophy, 2023; CFA Institute, 2024; American Bar Association, 2023; American Association for Public Opinion Research, 2021).
@@
Evidence Coverage aligns with evidence-provision requirements that evaluate whether claims are accompanied by verifiable supporting sources in accordance with established evidentiary and reporting guidelines (Bastian and Moher, 2021; Data Citation Synthesis Group, 2014; Chang et al., 2024; International Society for Stem Cell Research, 2025; CFA Institute, 2020). Claim Amount corresponds to content-adequacy criteria that assess whether a report supplies all necessary factual and contextual material needed to justify its reasoning under recognized completeness and transparency standards (Bastian and Moher, 2021; International Organization for Standardization, 2011; EQUATOR Network, 2025; Linguistic Society of America, 2024; The Journal of Organic Chemistry, 2025; U.S. Geological Survey, 2024; Global Reporting Initiative, 2023). Citation Amount align with source-attribution rules that examine whether in-text citations are provided at appropriate argumentative locations following authoritative norms for evidentiary traceability (International Committee of Medical Journal Editors, 2025; Wilkinson et al., 2016; IEEE, 2025; American Chemical Society, 2021; Harvard Law Review Association, 2020). Reference Amount are evaluated to determine whether the number of actually used references is appropriate and whether they collectively form a well-documented, traceable, and accessible evidence base (Stanford HAI, 2025; Data Citation Synthesis Group, 2014; American Political Science Association, 2012b).
@@
Claim Factuality reflects evidence-verification provisions that assess whether factual assertions in a report correspond to verifiable sources and documented evidence traces within established evaluative frameworks (Wilkinson et al., 2016; American Physical Society, 2023; National Institute of Standards and Technology, 2023; Navas and others, 2024; National Institute of Standards and Technology, 2024; U.S. Securities and Exchange Commission, 2024). Citation Support corresponds to source-justification checks that evaluate whether cited references substantively support the claims they accompany according to recognized standards for evidential accountability (Wilkinson et al., 2016; American Sociological Association, 2018; American Historical Association, 2024; Berez-Kroeker and others, 2018). Reference Support aligns with source-quality and provenance requirements that examine whether referenced materials reliably and transparently support the arguments for which they are cited, in line with authoritative data-governance and reporting criteria (Wilkinson et al., 2016; Chang et al., 2024; Data Citation Synthesis Group, 2014; American Psychological Association, 2025; International Accounting Standards Board, 2024). Reference Quality evaluates whether the set of sources consists of reproducible, trustworthy, and methodologically sound information sources, ensuring that the report’s evidentiary foundation is grounded in reliable materials (American Chemical Society, 2021; IEEE, 2025; Stanford Encyclopedia of Philosophy, 2025; World Medical Association, 2013; National Commission for the Protection of Human Subjects of Biomedical and Behavioral Research, 1979). Reference Diversity evaluates whether the evidence base avoids excessive concentration on a small number of sources and instead maintains a balanced level of source diversity, reducing the risk of narrow, biased, or selectively framed arguments (EQUATOR Network, 2025; Linguistic Society of America, 2024; American Association for Public Opinion Research, 2021).
Overall, this framework comprehensively integrates authoritative standards and guidelines from a wide spectrum of disciplines, ranging from quantitative sciences and engineering to the humanities, social sciences, and professional fields such as law and finance (see Table 10). By synthesizing the core principles shared across these diverse domains, this taxonomy prioritizes criteria that are universally recognized as essential for high-quality intellectual work. Consequently, the resulting evaluation model ensures that the generated reports not only adhere to domain-specific best practices but also satisfy the fundamental requirements of validity, clarity, and professional integrity demanded by the global research and practitioner communities.
@@
Appendix D Rubric Structure
In this appendix, we summarize the rubric taxonomy and scoring procedure used to evaluate Deep Research reports.
@@ Appendix D Rubric Structure
1. Request Fulfillment └ 1.1 Completeness 1.1.1 Criterion Does the report include all required elements without omission and present each clearly? [1.1.1.1 (Coverage)] The report must include all elements required by the User Query and the Expert Evaluation Guidance (EG), and each element must be presented with clear and understandable explanations. Completeness is judged against the EG; if the explanation for any required element falls short of the EG standard, that element is considered omitted. [1.1.1.2 (Coverage)] Each major requirement from the User Query must be developed in at least one sufficiently substantial paragraph. This length must consist of explanations directly relevant to the User Query and the EG context; if the content is only filler without substantive relation, the requirement is considered unfulfilled. [1.1.1.3 (Quality)] For each required element, the report must be supported by appropriate evidence, reasoning, and validation, guided by the EG. [1.1.1.4 (Quality)] Each required element must be supported with sufficient evidence and depth.
2. Analytical Soundness └ 2.2 Reasoning 2.2.5 Criterion All major claims follow logically from the previously presented facts, data, interpretations, and assumptions, without skipped steps or unsupported leaps. [2.2.5.1 (Coverage)] All claims that require logical support are explicitly linked to the relevant facts, data, and interpretations (including the key evidence specified in the EG), and no central claim is left without the necessary supporting evidence. [2.2.5.2 (Quality)] Each major claim is explained through clear, well-structured reasoning that shows how the underlying facts, data, and interpretations lead to that conclusion, making the logical connection explicit and easy for an expert reader to follow.
3. Structural Coherence └ 3.1 Introduction 3.1.1 Criterion Does the introduction clearly present the report’s topic, problem, and significance, avoiding excessive generalization or irrelevant topic development? Does it also provide sufficient context and motivation for the reader? [3.1.1.1 (Coverage)] The introduction must include the report’s topic, problem, and significance, and provide sufficient background and motivation for the reader to understand the report’s context and rationale. [3.1.1.2 (Quality)] The introduction must be sufficiently developed for a professional report, and each component specified in 3.1.1.1 must be treated with adequate depth. [3.1.1.3 (Quality)] Each component must be described clearly and specifically, without excessive generalization or ambiguity. [3.1.1.4 (Quality)] The introduction must present its components in a logical, coherent flow so the reader can easily grasp the report’s overall direction.
4. Format & Style └ 4.2 Writing Quality 4.2.3 Criterion Are technical terms defined when they first appear and used consistently thereafter? [4.2.3.1 (Coverage)] Technical terms and field-specific concepts must be defined when they are central to the argument, potentially ambiguous, or not guaranteed to be known by the intended audience. Well-established terms that are standard in the field do not require formal definitions if their meaning is clear from context. [4.2.3.2 (Coverage)] After being defined, technical terms must be used consistently with that definition throughout the document, including abbreviations and symbols.
5. Ethics & Compliance └ 5.2 Safety & Impact 5.2.1 Criterion Are the potential impacts of proposed policies, technologies, strategies, or research outcomes sufficiently considered, including key implications, possible side-effects, and interpretations from multiple perspectives (when essential)? [5.2.1.1 (Coverage)] Potential side-effects or limitations are discussed. [5.2.1.2 (Coverage)] Multiple perspectives and relevant contextual considerations are included. [5.2.1.3 (Quality)] Key implications are presented in a balanced way, and relevant contexts are sufficiently considered. [5.2.1.4 (Quality)] Each identified impact is analyzed with adequate detail, supported by data, evidence, or clear reasoning.
Table 11: Example criteria and rubric items from a four-level taxonomy (dimension, sub-dimension, criterion, and item: Coverage or Quality), with one example shown for each report-quality dimension.
D.1 Hierarchical Levels
We describe the rubric in two parts: (i) a shared taxonomy of dimensions and sub-dimensions, and (ii) the scoring instantiation for report-quality assessment, which specifies criteria and rubric items.
@@ D.1 Hierarchical Levels
Each dimension is decomposed into finer sub-dimensions (2–5 per dimension), specifying which aspects to examine within that dimension.
D.2 Criteria and Rubric Items
For the report-quality dimensions, we further specify criteria and rubric items. For the information-verification dimensions, we instead use metric-based measurements at the sub-dimension level; see Appendix F.4.
Level 3: Criteria (46)
Criteria operationalize each sub-dimension into evaluation requirements that are directly assessable from the report. One or more criteria are defined under each sub-dimension, and each criterion corresponds to a concrete aspect of report content, such as "Inclusion of requested items" and "Specification of scope/limitations."
Level 4: Rubric Items (101)
Rubric items at the fourth level decompose each criterion (Level 3) into atomic scoring units under two aspects, Coverage (C) and Quality (Q). Each rubric item is labeled as either (Coverage) or (Quality), and rubric items serve as the minimum unit for scoring. Table 11 shows example criteria and their associated rubric items.
D.3 Coverage and Quality
At Level 4, rubric items evaluate each criterion (Level 3) from two perspectives: Coverage (C) and Quality (Q). Coverage assesses whether the criterion is covered without omission as required throughout the document, including requirements distributed across multiple locations or composed of multiple detailed components. In contrast, Quality assesses whether the written content exhibits sufficient depth, logic, and rigor under professional report standards, conditioning on what is actually written (i.e., “how well it is written”). By separating Coverage and Quality, we can independently measure (1) what is missing and (2) the quality of what is included in a long report. In our rubric, Level 4 consists of 101 rubric items in total, comprising 66 Coverage items and 35 Quality items. Both use a 1–10 point scale, and the interpretation of score bands is summarized in Table 12.
@@ D.3 Coverage and Quality
D.4 Scoring and Aggregation
The score aggregation flow in this evaluation framework consists of rubric item $\rightarrow$ criterion $\rightarrow$ sub-dimension $\rightarrow$ dimension. First, rubric items (the lowest-level units) are evaluated with a 1–10 score (or N/A). Coverage (C) and Quality (Q) are computed separately at the rubric-item level, then integrated at the criterion level, and the resulting criterion scores are aggregated to the sub-dimension and dimension levels. N/A items are excluded from all average calculations.
Let $r$ be the report, and type $T\in\{C,Q\}$ represent Coverage and Quality, respectively. Let the score of each rubric item $i$ be $s_{r,i}\in\{1,\dots,10\}$ (or N/A), and let the set of non-N/A rubric items of type $T$ belonging to criterion $c$ be $I_{c,r}^{T}$. Let $\mathcal{C}_{s,r}$ be the set of criteria belonging to sub-dimension $s$, and let $\mathcal{S}_{d,r}$ be the set of sub-dimensions belonging to dimension $d$. Then, the rubric item $\rightarrow$ criterion $\rightarrow$ sub-dimension $\rightarrow$ dimension aggregation is defined as follows:
The Coverage/Quality score (criterion level) of report $r$ for criterion $c$ is
$S_{r}^{T}(c)=\frac{1}{\lvert I_{c,r}^{T}\rvert}\sum_{i\in I_{c,r}^{T}}s_{r,i},\quad T\in\{C,Q\},$ (2)
and if $I_{c,r}^{T}=\varnothing$, $S_{r}^{T}(c)$ is set to N/A.
The integrated criterion score $S_{r}(c)$ is set as the average of defined values among C/Q. Let the set of defined types for criterion $c$ be $\mathcal{T}_{c}=\{T\in\{C,Q\}\mid S_{r}^{T}(c)\neq\text{N/A}\}$.
$S_{r}(c)=\frac{1}{\lvert\mathcal{T}_{c}\rvert}\sum_{T\in\mathcal{T}_{c}}S_{r}^{T}(c).$ (3)
If $\mathcal{T}_{c}=\varnothing$, $S_{r}(c)$ is set to N/A.
The score of sub-dimension $s$ is defined as the average of criterion scores belonging to that sub-dimension.
$S_{r}(s)=\frac{1}{\lvert\mathcal{C}_{s,r}\rvert}\sum_{c\in\mathcal{C}_{s,r}}S_{r}(c).$ (4)
If $\mathcal{C}_{s,r}=\varnothing$, $S_{r}(s)$ is set to N/A.
The score of dimension $d$ is defined as the average of sub-dimension scores belonging to that dimension.
$S_{r}(d)=\frac{1}{\lvert\mathcal{S}_{d,r}\rvert}\sum_{s\in\mathcal{S}_{d,r}}S_{r}(s).$ (5)
If $\mathcal{S}_{d,r}=\varnothing$, $S_{r}(d)$ is set to N/A.
In summary, we compute $S_{r}^{C}(c),S_{r}^{Q}(c)$ by averaging rubric items within each criterion, integrate them to obtain the criterion score $S_{r}(c)$, then average criterion scores to obtain the sub-dimension score $S_{r}(s)$, and finally average sub-dimension scores to obtain the dimension score $S_{r}(d)$.
Appendix E Human-based Information Verification Protocol
E.1 Overview
@@
E.2 Step 1: Claim Extraction and Classification
@@
Evaluators segmented the report into paragraphs and sentences (in ‘Lx.Sy’ format, where L denotes the paragraph index and S denotes the sentence index), reviewed each sentence individually to identify sentences containing claims, and extracted only the core claims. Pronouns were replaced with explicit references according to the context, and if a single sentence contained multiple claims, they were separated. All claims were classified into 6 types (A–F): Explicit Citation (A), Implicit – Same Section (B), Implicit – Previous Section (C), Structural Recap (D), No Citation Required (E), and Unknown Source (F). For A–C type claims, the corresponding citation or evidence position was recorded together. This process was performed on 2 randomly selected reports (total 728 claims), thereby constructing a Ground Truth for evaluating the recall of the extraction model.
@@
E.3 Step 2: Claim Verification